# MELTing point: Mobile Evaluation of Language Transformers

### v2

@@ 5.3. Micro Experiments & Bottlenecks
(a) Temperature after one full conversation on Zephyr-3B (4-bit) on MLC-LLM
(b) Memory trace when running Zephyr-3B (4-bit) on LLMFarm
@@ 5.3.1. ML Operations
We start by introspecting Llama-7B (3-bit) on Android. We compile a custom version of TVM and MLC-LLM where we enable the vm_profiler in the backend and report kernel runtimes per operator of interest. In this section, we only measure per kernel latency, as the end-to-end latency is heavily impacted by the use of the profiler. Results are shown in Fig. 8 for the prefill, embed and decode operations. We see that most of the execution is taken up by de-quantize and matrix multiplication fused operations for the prefill and decode operations, taking up 97% and 95.7% of the total runtime, respectively. We hypothesize that the dequantization operation is also why 3-bit quantized networks may have performed worse than their 4-bit counterparts, as we discussed in Sec. 5.2.2. On the contrary, the embed operation seems mostly to be doing tensor conversion and retrieval operations. Since the generation process is mostly bottlenecked by the decode operation (evident also in Fig. 3 and 7), we proceed to investigate the real system bottleneck during execution via profiling. Due to lack of GPU tracing via the Android GPU Inspector on Galaxy S23, we apply the analysis on the iPhone 14 Pro.
@@ 5.5. Runtime at the Edge
Results for both scenarios are presented in Fig. 11. Specifically, in Fig. 10(a) we show the generation throughput (in tokens/sec) of various models on different Jetson devices and energy profiles, as run with llama.cpp on CUDA. We see that throughputs largely follow a monotonic trajectory with respect to model size and energy modes, with the notable exception of Orin Nano and Orin AGX at 15W, with the former performing +7.89 tokens/sec better on average. Overall, generation throughput is significantly higher than the equivalent mobile runtime, and this runtime can also be sustained for longer periods, as shown in Fig. 7. Indicatively, for Zephyr-7B (4-bit), the average throughput is 3.3$\times$ and 1.78$\times$ higher, for prefill and generation respectively.
@@ 7. Related Work
Edge execution of LLMs. There have been various lines of work attempting to port LLM computation on-device. Starting with frameworks, llama.cpp (Gerganov, 2023) and MLC (team, 2023) have stood out, offering cross-platform accelerated execution and support for various LLM architectures and device targets. Other open-source frameworks include llama2.c (Karpathy, 2023), aimed at simplicity without dependencies and tinygrad (tinygrad, 2023), focused on accelerated execution, but without support quantized mobile execution. Last, TinyChatEngine (mit-han lab, 2023) showcased on-device inference with compressed models, but lacks mobile support. Lately, OS providers have released their own platforms, such as Apple’s MLX (Hannun et al., 2023) and Google’s AICore (android.com, 2023). The former only provides support for desktop platforms (M-series SoCs) and the latter remains closed-source and only deployed on Pixel 8 Pro. Very recently, Google also released MediaPipe (Mark Sherwood, 2024) for running LLMs on device.

### v4

@@ 5.3. Micro Experiments & Bottlenecks
(a) Temperature after a full conversation on Zephyr-3B (4-bit) on MLC-LLM
(b) Memory trace when running Zephyr-3B (4-bit) on LLMFarm (GPU)
@@ 5.3.1. ML Operations
We start by introspecting Llama-7B (3-bit) on Android. We compile a custom version of TVM and MLC-LLM where we enable the vm_profiler in the backend and report kernel runtimes per operator of interest. In this section, we only measure per kernel latency, as the end-to-end latency is heavily impacted by the use of the profiler. Results are shown in Fig. 9 for the prefill, embed and decode operations. We see that most of the execution is taken up by de-quantize and matrix multiplication fused operations for the prefill and decode operations, taking up 97% and 95.7% of the total runtime, respectively. We hypothesize that the dequantization operation is also why 3-bit quantized networks may have performed worse than their 4-bit counterparts, as we discussed in Sec. 5.2.2. On the contrary, the embed operation seems mostly to be doing tensor conversion and retrieval operations. Since the generation process is mostly bottlenecked by the decode operation (evident also in Fig. 4 and 8), we proceed to investigate the real system bottleneck during execution via profiling. Due to lack of GPU tracing via the Android GPU Inspector on Galaxy S23, we apply the analysis on the iPhone 14 Pro.
@@ 5.5. Runtime at the Edge
Results for both scenarios are presented in Fig. 12. Specifically, in Fig. 12(a) we show the generation throughput (in tokens/sec) of various models on different Jetson devices and energy profiles, as run with llama.cpp on CUDA. We see that throughputs largely follow a monotonic trajectory with respect to model size and energy modes, with the notable exception of Orin Nano and Orin AGX at 15W, with the former performing +7.89 tokens/sec better on average. Overall, generation throughput is significantly higher than the equivalent mobile runtime, and this runtime can also be sustained for longer periods, as shown in Fig. 8. Indicatively, for Zephyr-7B (4-bit), the average throughput is 3.3$\times$ and 1.78$\times$ higher, for prefill and generation respectively. CPU runtimes are also provided in Appendix A.2.
@@ 7. Related Work
Edge execution of LLMs. There have been various lines of work attempting to port LLM computation on-device. Starting with frameworks, llama.cpp (Gerganov, 2023) and MLC (team, 2023) have stood out, offering cross-platform accelerated execution and support for various LLM architectures and device targets. Other open-source frameworks include llama2.c (Karpathy, 2023), aimed at simplicity without dependencies and tinygrad (tinygrad, 2023), focused on accelerated execution, but without support quantized mobile execution. Last, TinyChatEngine (mit-han lab, 2023) showcased on-device inference with compressed models, but lacks mobile support. Lately, OS providers have released their own platforms, such as Apple’s MLX (Hannun et al., 2023) and Google’s AICore (android.com, 2023). The former only provided support for desktop platforms (M-series SoCs) at the time of writing and the latter remains closed-source and only deployed on Pixel 8 Pro. Very recently, Google released MediaPipe (Mark Sherwood, 2024) for on-device LLM execution.
@@ 8. Conclusion
Supplementary Material
Appendix A Additional Evaluation Results
A.1. GPU Runtime of LLMFarm on iOS Devices
Table 5. LLMFarm (iOS) GPU performance and discharge rate for various models on mid-tier devices
Device Model Quantization Throughput (tokens/sec) Discharge per Token (mAh/token)
iPhone 14 Pro TinyLlama-1.1B q3_k 22.9826±1.0557 0.0085±0.0010
TinyLlama-1.1B q4_k 24.6531±1.1295 0.0079±0.0035
Gemma-2B q3_k 20.7053±0.0949 0.0378±0.0091
Gemma-2B q4_k 23.7938±0.1190 0.0322±0.0101
Zephyr-3B q3_k 10.5899±0.2858 0.0130±0.0088
Zephyr-3B q4_k 14.8165±0.4524 0.0090±0.0011
Llama2-7B q3_k 5.9889±0.1374 0.0528±0.0105
iPhone SE TinyLlama-1.1B q3_k 30.5524±1.1284 0.0074±0.0002
TinyLlama-1.1B q4_k 31.3951±1.0936 0.0066±0.0001
Gemma-2B q3_k 16.6111±0.0713 0.0296±0.0004
Gemma-2B q4_k 16.7310±0.1718 0.0289±0.0007
Zephyr-3B q3_k 13.7265±0.2955 0.0263±0.0002
Zephyr-3B q4_k 12.1618±1.4816 0.0364±0.0027
A.2. CPU Runtime of llama.cpp on Jetson Devices
Table 6. Jetson CPU performance and energy per token for various models and energy profiles.
Device Model Quantization Throughput (tokens/sec) Energy per token (mWh/token)
Orin AGX @ 50W TinyLlama-1.1B q4_k 13.3085±0.7917 0.0015±0.0004
Gemma-2B q4_k 6.2280±0.1455 0.0063±0.0006
Zephyr-3B q4_k 5.4001±0.2857 0.0033±0.0013
Mistral-7B q4_k 2.2248±0.0748 0.0201±0.0033
Llama2-7B q4_k 2.3284±0.0875 0.0204±0.0035
Orin AGX @30W TinyLlama-1.1B q4_k 10.7740±0.6574 0.0023±0.0007
Gemma-2B q4_k 4.8950±0.0925 0.0092±0.0016
Zephyr-3B q4_k 4.2830±0.2189 0.0067±0.0054
Mistral-7B q4_k 1.7442±0.0524 0.0181±0.0056
Llama2-7B q4_k 1.8100±0.0644 0.0102±0.0043